• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name eschulte

    Log entry time 17:11:50 on October 17, 2006

    Entry number 187416

    keyword=CHL Flow & the Cryo Target

    The 4K Flow from CHL dropped again around 16:00, warming the target up to about 25 degrees C. We called J. P. an asked him to contact the cryo group, as this seems to be becoming a late afternoon habbit. It turned out that to keep the flow rate high enough to prevent overpressure in the CHL return lines, the cryo group fed some 4K to the ESR. The cryo group, after talking with the target group, removed the extra flow to ESR, since the target group reported that we were in a stable running condition. The hall was not informed of this operation. As a result of the removal of the extra coolant flow, the pressure problem returned and the coolant flow from the CHL was valved off. This, of course caused the target to warm up.

    It has been decided, then that we should run not less than 7 g/s coolant flow, keeping the high power heaters around 350 to 400 W with the beam off. In the event that we loose coolant again, target operators are to do nothing until the situation stablizes itself, then establish running conditions.

    We are presently working to get the target back.